OpenStudy (anonymous):

Need help please! The club usually starts fundraising at the beginning of the year to ensure it raises enough money. Ethan tells the members that anyone who is planning to go on the trip will need to sell 10 raffle tickets and 5 admission tickets at the annual fair. There are 28 members who plan to go on the trip. (b) than says they will need to collect $3000 for the trip. Each admission ticket is being sold for $5. How much will they need to sell each raffle ticket for to be sure they have at least $3000? Ethan wants the tickets to be sold for a whole-dollar amount. Show your work.

OpenStudy (anonymous):

Please read, it's not too difficult, it's just a lot of words

OpenStudy (anonymous):

My guess is\[700\times r \ge3000\] which got me 4, that doesn't sound right ,so I want my answer to get checked

OpenStudy (anonymous):

@Owlcoffee

OpenStudy (anonymous):

@mathstudent55

OpenStudy (dumbcow):

the equation would look like this \[28(10r + 5(5)) \ge 3000\] \[280r \ge 3000 - 700\] \[r \ge \frac{2300}{280}\]

OpenStudy (anonymous):

Oh, thank you so much! would it come out as 8 ?

OpenStudy (dumbcow):

no you would need to round up... if its 8 then total money will be less than 3000

OpenStudy (anonymous):

And how?

OpenStudy (dumbcow):

?? what is 2300/280

OpenStudy (anonymous):

Smartcow, does that mean division? @dumbcow

OpenStudy (anonymous):

@Spicytaco1 what does he/she mean?

OpenStudy (spicytaco1):

yes divison

OpenStudy (anonymous):

I did division and got 8 when rounded

OpenStudy (dumbcow):

what did you get before rounding

OpenStudy (anonymous):

8.214

OpenStudy (dumbcow):

ok \[r \ge 8.214\] since r must be greater, then you must go up to 9, not down to 8

OpenStudy (anonymous):

so the answers is 9?